Four men and three women apply for a job at a company. three of the applicants are chosen for interviews. Let X denote the number of women in the interview pool. find the CDF of X.
this is what i did so far...
there is 35 possibilities,
P(0)= P(chosen 0 women)= (4 C 3) /35 = 4/35
P(1)= P(1 women)
WMM = (4 C 2)(3 C 1)/35 = 6*3 / 35 = 18/35
P(2)= P(2 women)
WWM = (4 C 1)(3 C 2)/35 = 4*3 / 35 = 12/35
P(3)= P(3 women)= WWW=(3 C 3) /35 = 1/35
BUT IM NOT SURE WHAT TO DO NEXT...

Answer by stanbon(75887)

Four men and three women apply for a job at a company. three of the applicants are chosen for interviews. Let X denote the number of women in the interview pool. find the CDF of X.
this is what i did so far...
there is 35 possibilities,
P(0)= P(chosen 0 women)= (4 C 3) /35 = 4/35
P(1)= P(1 women)
WMM = (4 C 2)(3 C 1)/35 = 6*3 / 35 = 18/35
P(2)= P(2 women)
WWM = (4 C 1)(3 C 2)/35 = 4*3 / 35 = 12/35
P(3)= P(3 women)= WWW=(3 C 3) /35 = 1/35
BUT IM NOT SURE WHAT TO DO NEXT...
-------------------
That's it. You have done what you were asked to do.
Cheers,
Stan H.